### v2

# The Beilinson–Bloch conjecture for some non-isotrivial varieties over global function fields

Abstract.
The Beilinson–Bloch conjecture is a generalization of the Birch and Swinnerton-Dyer conjecture, which relates the ranks of Chow groups of smooth projective varieties over global fields to the order of vanishing of $L$-functions. We prove the conjecture for certain classes of non-isotrivial varieties over ${\mathbb{F}}_{q}(t)$, including some cubic threefolds. We deduce the Birch and Swinnerton-Dyer conjecture for their intermediate Jacobians, and use it to establish new cases of the Tate conjecture over finite fields. We also prove some additional results on the arithmetic of these intermediate Jacobians.
1. Introduction
Let $X$ be a smooth projective variety over a global field $k$. The Beilinson–Bloch conjecture ([Bloch1984], [Beilinson1987]) predicts that the rank of the homologically trivial part of the codimension-$i$ Chow group of $X$ is equal to the order of vanishing at $s=i$ of the Hasse–Weil $L$-function $L(H^{2i-1}(X_{\overline{k}},{\mathbb{Q}}_{\ell}),s)$. For $X$ an abelian variety and $i=1$, this is equivalent to the (weak) Birch and Swinnerton-Dyer conjecture for $X$. In general, few cases of the Beilinson–Bloch conjecture have been established unconditionally, particularly over number fields, where the $L$-function is not even known in general to be well-defined at $s=i$. We recall some existing results on the conjecture in subsection 4.2.
In this paper we will be concerned with the Beilinson–Bloch conjecture in positive characteristic. Tate [Tate1965] and Artin–Tate [Tate1966b] recognized in the 1960s that the Birch and Swinnerton-Dyer conjecture over global function fields was closely related to the Tate conjecture for divisors on varieties over finite fields. They conjectured a special value formula for the zeta function of a smooth projective surface over a finite field, and predicted that under mild hypotheses, when the surface was fibered over a curve, the validity of that formula was equivalent to the strong Birch and Swinnerton–Dyer conjecture for the Jacobian of the generic fiber. This equivalence, termed Artin–Tate’s conjecture (d), was proved in the early 2000s by combining results of Artin–Tate, Grothendieck [Grothendieck1968], Milne [Milne1986], and Kato–Trihan [KT2003]. See [LRS2022] for some additional historical background, and two alternative proofs of conjecture (d).
Geisser in [Geisser2021, Theorem 1.1] established a higher-dimensional generalization of conjecture (d), which applies (under mild hypotheses) to any smooth proper variety $X$ over a finite field which is fibered over a curve. Geisser’s theorem can be reformulated as saying that the Tate conjecture for divisors on $X$ is equivalent to the Tate conjecture for divisors on the generic fiber, together with the Birch and Swinnerton-Dyer conjecture for the Albanese variety of the generic fiber. In [Broe2025, Corollary 5.6], we generalized part of Geisser’s result to cycles of higher codimension, giving a sufficient condition for the Beilinson–Bloch conjecture to hold for a variety over a function field. From this we deduced some new cases of the conjecture for isotrivial varieties (similar isotrivial examples were known previously, such as those given in [Jannsen2007, Remark 5.2] and [Kahn2021, Corollary 1]). Here a variety over a global function field $K/{\mathbb{F}}_{q}$ is called isotrivial if its base change to $\overline{K}$ is defined over $\overline{{\mathbb{F}}}_{q}$.
It remains to show that this method can prove cases of the Beilinson–Bloch conjecture in higher codimension for interesting non-isotrivial varieties. Since our criterion for the conjecture requires some control over the bad reduction of a variety over a function field, it is natural to seek examples where the singularities of the special fibers are as mild as possible. We are thus led to consider Lefschetz pencils. Recall that a Lefschetz pencil for a smooth projective variety $X$ is roughly a generically smooth family of hyperplane sections of $X$, parametrized by $\mathbb{P}^{1}$, such that each singular member of the family (over the algebraic closure) is smooth away from one ordinary double point singularity. They have long played a structural role in algebraic and arithmetic geometry, notably in Deligne’s papers on the Weil conjectures ([Deligne1974], [Deligne1980]).
Theorem 1.1 (Theorem 6.4).
Let $X$ be a smooth cubic fourfold over ${\mathbb{F}}_{q}$, and let $f:Y\to\mathbb{P}^{1}_{{\mathbb{F}}_{q}}$ be a Lefschetz pencil for $X$. Then the Beilinson–Bloch conjecture holds for the generic fiber of $f$.
This establishes the conjecture for a class of smooth cubic threefolds over ${\mathbb{F}}_{q}(t)$. Under some stronger hypotheses on $X$, we use results of Jannsen [Jannsen2007] to compute all motivic cohomology groups (with rational coefficients) of the generic fiber of $f$, in terms of certain simpler invariants related to étale cohomology (theorem 6.12). We also prove cases of the Beilinson–Bloch conjecture for cubic hypersurfaces of higher dimension (theorem 6.5), and establish the finiteness of the Albanese kernel for a class of surfaces over ${\mathbb{F}}_{q}(t)$ (corollary 6.9). The varieties covered by these results are shown to be non-isotrivial in many cases (remark 6.11): for instance, the cubic threefolds of theorem 1.1 are never isotrivial.
The proof of theorem 1.1 relies on the Tate conjecture for cubic fourfolds over finite fields, established unconditionally in [Charles2013], [Madapusi2015], and [AKPW2025]. Note that a similar Lefschetz pencil argument was used by Charles–Pirutka [CP2015] in their proof of the integral Tate conjecture for smooth cubic fourfolds in characteristic $\neq 2,3$.
Smooth cubic threefolds over $\mathbb{C}$ have long been studied through their intermediate Jacobians, which are principally polarized abelian fivefolds constructed from the Hodge structures of the threefolds. Clemens–Griffiths [CG1972] famously took this approach to show that smooth complex cubic threefolds are never rational, a theorem soon extended by Murre [Murre1973] to base fields of characteristic not 2. Later work of Murre [Murre1985], or rather its generalization by Achter–Casalaina-Martin–Vial [ACMV2023], permits a definition of the intermediate Jacobian of a smooth cubic threefold $X$ over an arbitrary field. This abelian variety satisfies a certain universal property, and admits a concrete geometric interpretation: it is isomorphic to the Albanese variety of the Fano surface of lines on $X$ [Ciurca2024, Proposition 5.54]. Ciurca in loc. cit. has also announced a proof of the irrationality of smooth cubic threefolds in characteristic 2: his manuscript is currently only available as a preprint, and we do not rely on it for any of our results here.
There is a certain relation between the Chow motive of a smooth cubic threefold over any field and that of its intermediate Jacobian (proposition 3.4). It formally follows that the Beilinson–Bloch conjecture for a smooth cubic threefold over a global field is equivalent to the Birch and Swinnerton-Dyer conjecture for its intermediate Jacobian. From this and [KT2003], we deduce the strong form of BSD for the intermediate Jacobians of the cubic threefolds treated in theorem 1.1. Since the rational $\ell$-adic Tate modules of these abelian varieties are controlled by the well-understood monodromy of Lefschetz pencils, we can make several other observations about their arithmetic. In particular, they have everywhere semistable reduction, and their geometric isogeny classes are simple and non-isotrivial (proposition 7.1). We moreover explicitly compute their $L$-functions, and give an example of Mordell–Weil rank 28. In future work, we aim to undertake a more focused study of their BSD invariants.
We conclude the paper by proving some new cases of the Tate conjecture over finite fields (remark 7.4). Specifically, we consider a smooth cubic threefold $X$ over a global function field of characteristic $\geq 3$, and assume that $X$ satisfies the Beilinson–Bloch conjecture (which holds unconditionally for the cubic threefolds of theorem 1.1). Using the Tate conjecture for the Fano surface $F$ of $X$, proved by Roulleau [Roulleau2013], and the Birch and Swinnerton-Dyer conjecture for the intermediate Jacobian of $X$, combined with the backwards direction of the aforementioned theorem of Geisser, we establish the Tate conjecture for a certain auxiliary threefold over the constant field, obtained by spreading out $F$. It is interesting that here the shortest path between two cases of the Tate conjecture over finite fields seems to pass through the domain of function fields, and we hope that the argument demonstrates some utility of the general Beilinson–Bloch conjecture in positive characteristic.
Structure of the paper
Section 2 recalls some results on the structure of Chow groups of certain varieties, particularly cubic hypersurfaces. Section 3 introduces tools from the theory of Chow motives, and uses them to relate the motive of a smooth cubic threefold with that of its intermediate Jacobian. In section 4 we state Jannsen’s generalized Tate conjecture and the Beilinson–Bloch conjecture, and give sufficient conditions for them. Section 5 collects results from SGA 7 [SGA7-II] and Weil II [Deligne1980] on the monodromy of Lefschetz pencils. These permit an explicit computation (under mild hypotheses) of the $L$-function associated with the middle cohomology of the generic fiber of a Lefschetz pencil over ${\mathbb{F}}_{q}$, in terms of the zeta function of the total space of the pencil. Section 6 proves cases of the Beilinson–Bloch conjecture for generic fibers of certain Lefschetz pencils, essentially via geometric arguments. Section 7 deduces BSD for the intermediate Jacobian of a smooth cubic threefold over ${\mathbb{F}}_{q}(t)$ which arises from a Lefschetz pencil of a cubic fourfold, and uses it to obtain cases of the Tate conjecture over ${\mathbb{F}}_{q}$. Here we also employ the earlier results on monodromy of Lefschetz pencils to describe the arithmetic of the intermediate Jacobian, and compute its $L$-function.
Notation and conventions
The letter $k$ will denote a field (arbitrary unless stated otherwise), and $\overline{k}$ will denote a choice of algebraic closure. We let $G_{k}$ denote the absolute Galois group of $k$. A variety over $k$ is a finite-type separated integral $k$-scheme. By $\ell$ we will always mean a fixed prime not equal to the characteristic of the base field. Our conventions for motives follow André’s book [Andre2004], and also match those of the Stacks project [stacks-project, Tag 0FG9]. By default, all Chow groups and motives have rational coefficients, and are taken with respect to rational equivalence. For a smooth projective variety $X$, the Chow group modulo homological equivalence (with respect to $\ell$-adic cohomology) is denoted by $\operatorname{CH}^{i}(X)_{\hom}$, and the homologically trivial subgroup of $\operatorname{CH}^{i}(X)$ is denoted by $\operatorname{CH}^{i}(X)_{0}$. By $\operatorname{CH}_{i}(X)$ we mean the Chow group graded by dimension of cycles, rather than codimension. If $R$ is a commutative ring, the Chow group with coefficients in $R$ is denoted by $\operatorname{CH}^{i}(X)_{R}$. The category of motives over $k$ is denoted by $\mathcal{M}(k)$. The tensor unit in this category is denoted by $\mathbf{1}$. The motive of a smooth projective scheme $Y$ is denoted by $\mathfrak{h}(Y)$. The dual of a motive $M$ is denoted by $M^{\vee}$, and the dual of a morphism $f:M\to N$ of motives is denoted by $f^{\vee}:N^{\vee}\to M^{\vee}$. The $n$-fold Tate twist of $f$ is denoted by $f(n):M(n)\to N(n)$. Our sign convention for Tate twists of motives is such that $\mathfrak{h}(\mathbb{P}^{1})\cong\mathbf{1}\oplus\mathbf{1}(-1)$. Subscripts on schemes, motives, or morphisms thereof denote base change. By $H^{i}(Y_{\overline{k}},{\mathbb{Q}}_{\ell})$ we mean $\ell$-adic étale cohomology. For a finite field ${\mathbb{F}}_{q}$, $\operatorname{Frob}_{q}$ denotes the geometric Frobenius in $G_{{\mathbb{F}}_{q}}$, which sends $x\mapsto x^{1/q}$.
Acknowledgments
I thank Jeff Achter, Niranjan Ramachandran, Padma Srinivasan and Doug Ulmer for helpful conversations.
2. Preliminaries on algebraic cycles
We begin by collecting some known results on the Chow groups of certain varieties. Our focus is on cubic hypersurfaces (whose Chow groups are also surveyed in [Huybrechts2023, Section 7.7.4]). Throughout the paper we will freely use the following standard lemma.
Lemma 2.1.
For any finite-type scheme $X$ over $k$ and any algebraic field extension $L/k$, the pullback map $\operatorname{CH}_{i}(X)\to\operatorname{CH}_{i}(X_{L})$ is injective. When $X$ is additionally smooth, the pullback is also injective when $L/k$ is an arbitrary extension.
Proof.
The former claim is [stacks-project, Tag 0FH8]. The latter follows from the argument of [LS2025, Lemma 2.1]. This uses the refined Gysin homomorphism of [Fulton1998, Chapter 6.2]; since a closed immersion of finite-type smooth schemes is a regular embedding [Fulton1998, Appendix B.7.2], the Gysin homomorphism is well-defined in the generality needed here. ∎
Let $\Omega$ be a universal domain containing $k$, i.e. $\Omega$ is an algebraically closed field of infinite transcendence degree over its prime field.
Proposition 2.2 ([FV2023, Lemma 5.1]).
Let $X$ be a smooth projective variety over $k$, such that there exists a curve $C\subseteq X_{\Omega}$ whose points generate $\operatorname{CH}_{0}(X_{\Omega})$. If $H^{3}(X_{\overline{k}},{\mathbb{Q}}_{\ell})=0$, then rational and numerical equivalence agree on $\operatorname{CH}^{2}(X)$.
Remark 2.3.
In particular, the proposition applies when $X$ is a smooth cubic hypersurface of dimension $\geq 2$ and $\neq 3$: then $X_{\Omega}$ is unirational [Kollar2002, Theorem 1], hence rationally chain connected [Fulton1998, Example 10.1.6], so $\operatorname{CH}_{0}(X_{\Omega})_{\mathbb{Z}}\cong{\mathbb{Z}}$.
For a smooth, projective, geometrically connected variety $X$ over $k$, there exists a natural “Albanese” map from the group $\operatorname{CH}_{0}(X)_{{\mathbb{Z}},0}$ of zero-cycles of degree zero to the $k$-points of the Albanese variety $\operatorname{Alb}(X)$. We denote by $T(X)$ the kernel of this map, and refer to it as the Albanese kernel of $X$.
Proposition 2.4 ([KS1983, Proposition 9]).
When $k$ is finite, $T(X)$ is finite.
Corollary 2.5.
If $k$ is an algebraic extension of a finite field, then for any smooth cubic hypersurface $X$ over $k$ of dimension $\geq 3$, $\operatorname{CH}_{1}(X)\cong{\mathbb{Q}}$.
Proof.
It suffices to assume that $k=\overline{k}$. Then by [Shen2019, Theorem 1.7], there exists a surjective map $\operatorname{CH}_{0}(F(X))\to\operatorname{CH}_{1}(X)$. Here $F(X)$ is the Fano variety of lines on $X$, which is smooth and connected by [AK1977, Theorems 1.10 and 1.17]. Proposition 2.4 implies that $\operatorname{CH}_{0}(F(X))\cong{\mathbb{Q}}$, whence the result. ∎
We note in passing that [Shen2019, Theorem 1.7] has been generalized to cycles of higher dimension in [Lyu2025, Theorem 0.1], though we will not use this.
Work of Esnault–Levine–Viehweg gives some control over Chow groups of general projective schemes in a certain range.
Proposition 2.6 ([ELV1997, Theorem 4.6]).
Let $Z$ be a closed subscheme of $\mathbb{P}^{n}_{k}$, defined by $r$ homogeneous equations of respective degrees $d_{1}\geq...\geq d_{r}\geq 2$. If $j\in{\mathbb{Z}}$ and either $d_{1}\geq 3$ or $r\geq j+1$, then whenever
$\sum\limits_{i=1}^{r}\binom{j+d_{i}}{j+1}\leq n,$
we have $\operatorname{CH}_{j^{\prime}}(Z)\cong{\mathbb{Q}}$ for $0\leq j^{\prime}\leq j$.
For example, in the case of a cubic hypersurface $Z$ of dimension $\geq 5$ over any field, we find that $\operatorname{CH}_{1}(X)\cong{\mathbb{Q}}$.
Proposition 2.7.
Let $k$ be a finitely generated field and $X$ a smooth cubic fourfold over $k$. If $\operatorname{char}(k)\neq 2$, or $k$ is finite of characteristic 2, then the Tate conjecture holds for $X$.
Proof.
See [Charles2013, Corollary 6], [Madapusi2015, Theorem 5.14], and [AKPW2025, Theorem 4.9]. ∎
Remark 2.8.
We believe the Tate conjecture for smooth cubic fourfolds over finitely generated fields of characteristic 2 should follow from the case of a finite base field, via standard specialization arguments. However, this is beyond our present scope.
When $X$ is a smooth cubic hypersurface of dimension $d$, its middle Betti number $b_{d}(X)$ is computed in [Huybrechts2023, Corollary 1.1.12]: when $d=2,$ $3,$ or $4$, $b_{d}(X)$ is respectively 7, 10, or 23. We will freely use these values throughout the paper without further comment. We also note the classical result that a smooth cubic surface over an algebraically closed field is isomorphic to a blowup of $\mathbb{P}^{2}$ at six points [Huybrechts2023, Proposition 4.2.4].
3. Motives
In this section, we recall some needed background on motives, and prove a motivic relation between a smooth cubic threefold and its intermediate Jacobian (proposition 3.4). Some of our arguments will rest on the theory of finite-dimensional motives [Kimura2005, Definition 3.7], in which the nilpotence theorem plays a fundamental role.
Proposition 3.1 ([Kimura2005, Proposition 7.5]).
If $M$ is a finite-dimensional motive over $k$, and $f:M\to M$ is an endomorphism which induces the zero map modulo numerical equivalence, then $f$ is nilpotent.
If $M$ is a motive over $k$ and $L/k$ a field extension, then $M_{L}$ is finite-dimensional iff $M$ is [Vial2017, Theorem 2].
A motive of abelian type is an object of the thick and rigid subcategory of $\mathcal{M}(k)$ generated by the motives of abelian varieties. For instance, the motive of a smooth, projective, geometrically connected curve is of abelian type [Scholl1994, Proposition 4.5]. When $m\geq 1$ and $\operatorname{char}(k)\nmid m$, the Fermat hypersurface $X\subset\mathbb{P}^{n+1}_{k}$ defined by the equation
$\sum\limits_{i=0}^{n+1}x_{i}^{m}=0$
has motive of abelian type [KS79, Theorem 1]. Motives of abelian type are finite-dimensional [Kimura2005, Example 9.1].
Let $X$ be a smooth projective scheme over $k$ of pure dimension $d$, and let $H$ be a classical Weil cohomology theory on $\mathcal{M}(k)$ [Andre2004, 3.4].
Definition 3.2.
A Chow–Künneth decomposition of the motive of $X$ (with respect to $H$) consists of an isomorphism of motives
$\mathfrak{h}(X)\cong\bigoplus\limits_{0\leq i\leq 2d}\mathfrak{h}^{i}(X),$
such that for all $0\leq i\leq 2d$ and $j\in{\mathbb{Z}}$, $H^{j}(\mathfrak{h}^{i}(X))=0$ whenever $i\neq j$. A Chow–Künneth decomposition is said to be self-dual if for all $i$, the canonical isomorphism $\mathfrak{h}(X)^{\vee}\cong\mathfrak{h}(X)(d)$ induces an isomorphism $\mathfrak{h}^{i}(X)^{\vee}\cong\mathfrak{h}^{2d-i}(X)(d)$. Equivalently, if $\pi_{i}:\mathfrak{h}(X)\to\mathfrak{h}^{i}(X)$ denotes the projector onto the $i$-th Chow–Künneth component, the decomposition is self-dual iff for all $i$, $\pi_{i}$ is the transpose of $\pi_{2d-i}$.
By [Andre2004, Théorème 4.2.5.2 and Remarque 5.1.1.2], the definition does not depend on the choice of $H$, so we will simply speak of Chow–Künneth decompositions, with no reference to $H$. Self-dual Chow–Künneth decompositions are known to exist for motives of abelian varieties [DM1991], curves, surfaces [Murre1990], complete intersections in $\mathbb{P}^{n}$ [KS2016], and several other classes of varieties. All the concrete examples of Chow–Künneth decompositions considered in this paper will have the property that $\operatorname{CH}^{i}(\mathfrak{h}^{1}(X))=0$ when $i\neq 1$, and $\operatorname{CH}^{1}(\mathfrak{h}^{1}(X))=\operatorname{CH}^{1}(X)_{0}$ (see [Scholl1994, Theorem 4.4]).
Proposition 3.3.
A smooth complete intersection $X\subset\mathbb{P}^{r}_{k}$ of dimension $d$ has a Chow–Künneth decomposition of the form
$\mathfrak{h}(X)\cong\left[\bigoplus\limits_{\begin{subarray}{c}0\leq i\leq d\\ i\neq d/2\end{subarray}}\mathbf{1}(-i)\right]\oplus\mathfrak{h}^{d}(X).$
When $X$ is a smooth cubic hypersurface of dimension $\leq 3$, the motive of $X$ is finite-dimensional.
Proof.
The existence and form of the Chow–Künneth decomposition is proved in [KS2016, Section 3.5]. Suppose $X$ is a smooth cubic hypersurface of dimension $\leq 3$. Curves have finite-dimensional motive, so we may assume $\dim X\geq 2$. When $X$ has a rational point, $X$ is unirational by [Kollar2002, Theorem 1], and [Vial2017, Theorem 3.12 and Proposition 3.13] then imply that $\mathfrak{h}(X)$ is finite-dimensional. Theorem 2 of loc. cit. yields the same conclusion when $X$ does not have a rational point. ∎
The following proposition is likely known to experts, but we did not find a proof over an arbitrary field in the literature. It was proved by Shermenev over an algebraically closed field [Shermenev1970], via a different method. See [Huybrechts2023, Remark 5.3.17] for some relevant background and references.
Proposition 3.4.
Let $X$ be a smooth cubic threefold over $k$, equipped with a self-dual Chow–Künneth decomposition. There is a canonical isomorphism
$\mathfrak{h}^{3}(X)\cong\mathfrak{h}^{1}(A)(-1),$
where $A$ is the Albanese variety of the Fano surface $F$ of lines on $X$.
Proof.
Fix a self-dual Chow–Künneth decomposition
$\mathfrak{h}(F)\cong\mathbf{1}\oplus\mathfrak{h}^{1}(F)\oplus\mathfrak{h}^{2}(F)\oplus\mathfrak{h}^{1}(F)(-1)\oplus\mathbf{1}(-2)$
as in [Scholl1994, Theorem 4.4], so that $\mathfrak{h}^{1}(F)\cong\mathfrak{h}^{1}(A)$. Let $I$ be the incidence correspondence
$I=\{(L,x):x\in L\}\subset F\times_{k}X,$
and let
$p=[I]\in\operatorname{CH}^{2}(F\times_{k}X)=\operatorname{Hom}_{\mathcal{M}(k)}(\mathfrak{h}(F),\mathfrak{h}(X)).$
Then
$p^{\vee}$ $\in\operatorname{Hom}_{\mathcal{M}(k)}(\mathfrak{h}(X)^{\vee},\mathfrak{h}(F)^{\vee})$
$=\operatorname{Hom}_{\mathcal{M}(k)}(\mathfrak{h}(X)(3),\mathfrak{h}(F)(2))$
$=\operatorname{Hom}_{\mathcal{M}(k)}(\mathfrak{h}(X),\mathfrak{h}(F)(-1)).$
Let $\pi_{i}:\mathfrak{h}(X)\to\mathfrak{h}^{i}(X)$ (resp. $\psi_{j}:\mathfrak{h}(F)\to\mathfrak{h}^{j}(F)$) denote the projectors onto the Chow–Künneth components. By [CP2015, page 5], $p^{\vee}$ induces an isomorphism $H^{3}(Z_{\overline{k}},{\mathbb{Z}}_{\ell})\cong H^{1}(F_{\overline{k}},{\mathbb{Z}}_{\ell}(-1))$, so that
$g=(\psi_{1}(-1))\circ\left(p^{\vee}|_{\mathfrak{h}^{3}(X)}\right)\circ\pi_{3}\circ\left(p|_{\mathfrak{h}^{3}(F)}\right):\mathfrak{h}^{1}(F)(-1)\to\mathfrak{h}^{1}(F)(-1)$
induces an isomorphism on $\ell$-adic cohomology. Then [Scholl1994, Proposition 4.5] implies that $g$ is an isomorphism of motives modulo rational equivalence. Since $g$ factors through $\pi_{3}:\mathfrak{h}(X)\to\mathfrak{h}^{3}(X)$, we thus obtain a direct sum decomposition of the form
$\mathfrak{h}^{3}(X)\cong\mathfrak{h}^{1}(F)(-1)\oplus C.$
As $H^{3}(X_{\overline{k}},{\mathbb{Q}}_{\ell})$ and $H^{1}(F_{\overline{k}},{\mathbb{Q}}_{\ell}(-1))$ are isomorphic, the $\ell$-adic cohomology of $C$ is zero. By the finite-dimensionality of $\mathfrak{h}(X)$ and the nilpotence theorem (proposition 3.1), we must have $C=0$. ∎
Remark 3.5.
Let $X$ be a general smooth cubic fivefold over a subfield $k$ of $\mathbb{C}$, and $F$ the Fano surface of planes on $X$. Collino in [Collino1986] proved that the incidence correspondence induces an isomorphism from $\operatorname{Alb}(F_{\mathbb{C}})$ to the intermediate Jacobian of $X_{\mathbb{C}}$. In characteristic zero, smooth cubic fivefolds are known to always have finite-dimensional motive [Vial2013, Example 4.15]. Imitating the argument of proposition 3.4, one can thus show that $\mathfrak{h}^{5}(X)\cong\mathfrak{h}^{1}(F)(-2)$. See also [Vial2013, Theorem 4] for a variation on this statement which does not require $X$ to be general.
If $X$ is instead a smooth cubic fivefold over an arbitrary field $k$, in general we do not know finite-dimensionality of the motive of $X$. However, when the Fano surface of planes $F$ is smooth, the same argument shows that $\mathfrak{h}^{1}(F)(-2)$ and $\mathfrak{h}^{5}(X)$ at least have a certain summand in common, which is isomorphic to $\mathfrak{h}^{1}(B)(-2)$ for some abelian variety $B$. The $\ell$-adic cohomology of this summand is the image of $H^{1}(F_{\overline{k}},{\mathbb{Q}}_{\ell}(-2))$ inside $H^{5}(X_{\overline{k}},{\mathbb{Q}}_{\ell})$. When $X$ additionally lifts to a cubic fivefold in characteristic zero which is covered by Collino’s result, one can show via proper base change, as in [CP2019, page 5], that this summand is all of $\mathfrak{h}^{1}(F)(-2)$.
We end the section by recalling the blowup formula for motives.
Proposition 3.6 ([Hanamura2011, Corollary 3.5]).
If $X$ is a smooth projective variety over $k$, $Z\subseteq X$ is a smooth closed subvariety of codimension $c$, and $\widetilde{X}$ is the blowup of $X$ along $Z$, then there is a canonical isomorphism
$\mathfrak{h}(\widetilde{X})_{\mathbb{Z}}\cong\mathfrak{h}(X)_{\mathbb{Z}}\oplus\bigoplus\limits_{j=1}^{c-1}\mathfrak{h}(Z)(-j)_{\mathbb{Z}}$
in the category of motives with $\mathbb{Z}$}-coefficients over $k$. This induces isomorphisms
$\operatorname{CH}^{i}(\widetilde{X})_{\mathbb{Z}}\cong\operatorname{CH}^{i}(X)_{\mathbb{Z}}\oplus\bigoplus\limits_{j=1}^{c-1}\operatorname{CH}^{i-j}(Z)_{\mathbb{Z}}$
for all $i$.
4. Conjectures on algebraic cycles
We now review the conjectures on algebraic cycles we will be concerned with, and give some sufficient conditions for them.
4.1. The generalized conjectures of Tate and Beilinson
For $k$ a field, $X$ a scheme with finite-type, separated structure morphism $f:X\to\operatorname{Spec}k$, and $i,j\in{\mathbb{Z}}$, we define the $\ell$-adic étale homology of $X$ as in [Jannsen1990, Example 6.7], via
$H_{i}(X_{\overline{k}},{\mathbb{Q}}_{\ell}(j))=H^{-i}(X_{\overline{k}},(R{\overline{f}}^{!}{\mathbb{Q}}_{\ell})(-j)).$
Here $\overline{f}$ is the base change of $f$ to $\overline{k}$.
Define the (rational) motivic homology and cohomology of $X$ as in [Jannsen1990, 5.12 and Example 6.12], via
$H_{i}^{\mathcal{M}}(X,{\mathbb{Q}}(j))$ $=K^{\prime}_{i-2j}(X)^{(-j)},$
$H^{i}_{\mathcal{M}}(X,{\mathbb{Q}}(j))$ $=K_{2j-i}(X)^{(j)}.$
Here $K_{m}(X)$ denotes the $m$-th Quillen $K$-theory of vector bundles on $X$, and $K_{m}(X)^{(r)}$ denotes the simultaneous eigenspace of $K_{m}(X)\otimes_{\mathbb{Z}}{\mathbb{Q}}$ for the Adams operations $\{\psi^{n}:n\in{\mathbb{Z}}-\{0\}\}$, on which $\psi^{n}$ acts as multiplication by $n^{r}$. The definition of $K^{\prime}_{m}(X)^{(r)}$ is the same, except using the $K$-theory of coherent sheaves on $X$ instead of vector bundles. See [Soule1985, 1.5 and 7.2] for the definition of the Adams operations.
For context, we note that for a smooth variety $X$ over $k$, and $i$, $j\in{\mathbb{Z}}$, there is a a canonical isomorphism
$H^{i}_{\mathcal{M}}(X,{\mathbb{Q}}(j))\cong\operatorname{CH}^{j}(X,2j-i)_{\mathbb{Q}},$
cf. [Bloch1986, page 269(vii)] and [Jannsen1990, Example 6.12]. On the right is Bloch’s higher Chow group (with $\mathbb{Q}$}-coefficients). For such an $X$, we thus let $H^{i}_{\mathcal{M}}(X,{\mathbb{Z}}(j))$ denote $\operatorname{CH}^{j}(X,2j-i)_{\mathbb{Z}}$. The more modern definitions of integral motivic cohomology and Borel–Moore homology are not in terms of higher Chow groups or $K$-theory [MVW2006, Definitions 3.4 and 16.20], but this comparatively classical approach will suffice for us.
In any case, we will not be too preoccupied with the precise definition of motivic (co)homology. Both étale and motivic (co)homology form twisted Poincaré duality theories in the sense of [Jannsen1990, Definition 6.1] [^1] There is a caveat in the case of motivic (co)homology, which forms an “absolute” twisted Poincaré duality theory, rather than a relative one. See [Jannsen1990, Example 6.12]., so that we have the following purely formal tools for working with them. Out of an abundance of caution, in the remainder of this section we work in the category of quasi-projective schemes over the base field, unless otherwise noted [^2] We could possibly get away with weaker hypotheses, but [Soule1985] as stated works in this generality..
Proposition 4.1.
For $X$ a quasi-projective scheme over $k$, let $H_{i}(X,j)$ denote either rational motivic or étale homology, and $H^{i}(X,j)$ the corresponding cohomology.
  1. (i)
    The homology $H_{i}(X,j)$ is covariant in $X$ with respect to proper morphisms, and contravariant with respect to étale morphisms. The cohomology $H^{i}(X,j)$ is contravariant in $X$.
  2. (ii)
    For $0\leq i\leq\dim X$, there is a canonical isomorphism $H_{2i}^{\mathcal{M}}(X,{\mathbb{Q}}(i))\cong\operatorname{CH}_{i}(X)$ [Jannsen1990, Lemma 6.12.4]. For $i,j\in{\mathbb{Z}}$, there is a regulator map
    $H_{i}^{\mathcal{M}}(X,{\mathbb{Q}}(j))\otimes_{{\mathbb{Q}}}{\mathbb{Q}}_{\ell}\to H_{i}(X_{\overline{k}},{\mathbb{Q}}_{\ell}(j))$
    [Jannsen1990, Example 8.3].
  3. (iii)
    There is a cap product
    $\cap:H_{i}(X,j)\otimes H^{m}(X,n)\to H_{i-m}(X,j-n)$
    compatible with contravariance for étale morphisms. For a proper morphism $f:Y\to X$, the cap product satisfies the projection formula
    $f_{*}(\alpha\cap f^{*}(\beta))=f_{*}(\alpha)\cap\beta.$
  4. (iv)
    For $X$ irreducible of dimension $d$, there is a fundamental class $\eta_{X}\in H_{2d}(X,d)$, functorial with respect to contravariance for étale morphisms. When $X$ is additionally smooth, cap product with $\eta_{X}$ is an isomorphism
    $H^{2d-i}(X,d-j)\cong H_{i}(X,j)$ (4.1)
    compatible with regulator maps. Cohomology is thus covariant for proper morphisms of smooth varieties. These properties naturally extend to non-irreducible smooth schemes via
    $H^{i}(X\sqcup Y,j)\cong H^{i}(X,j)\oplus H^{i}(Y,j).$
    For $X$ smooth and irreducible, and $f:Y\to X$ finite étale of degree $m$, we have $f_{*}(\eta_{Y})=m\cdot\eta_{X}$, so the projection formula gives $f_{*}(f^{*}(\alpha))=m\cdot\alpha$ for any $\alpha\in H^{i}(X,j)$.
  5. (v)
    For $Z\subseteq X$ closed, with open complement U=XZU=X-Z, there is a long exact sequence
    $...\to H_{i}(Z,j)\to H_{i}(X,j)\to H_{i}(U,j)\to H_{i-1}(Z,j)\to$... (4.2)
    compatible with regulator maps.
In terms of motivic and étale homology, Jannsen generalized conjectures of Tate [Tate1994, Conjecture $T^{j}(X)$] and Beilinson [Beilinson1987, page 3] to arbitrary finite-type separated schemes over ${\mathbb{F}}_{q}$.
Conjecture 4.2 ($\operatorname{TB}_{i,j}(X)$: [Jannsen2007, Conjecture 4.14]).
For $X$ a finite-type separated scheme (not necessarily quasi-projective) over a finite field ${\mathbb{F}}_{q}$, and for $i,j\in{\mathbb{Z}}$,
  1. (i)
    the regulator map induces an isomorphism
    $H_{i}^{\mathcal{M}}(X,{\mathbb{Q}}(j))\otimes_{{\mathbb{Q}}}{\mathbb{Q}}_{\ell}\xrightarrow{\sim}H_{i}(X_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(j))^{G_{{\mathbb{F}}_{q}}};$
  2. (ii)
    the $\ell$-adic homology $H_{i}(X_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(j))$ is 1-semisimple.
Here we define a finite-dimensional $G_{{\mathbb{F}}_{q}}$-representation $V$ to be 1-semisimple if 1 is a semisimple eigenvalue of the geometric Frobenius $\operatorname{Frob}_{q}\in G_{{\mathbb{F}}_{q}}$ acting on $V$. We denote the truth value of the conjecture for a given $(X,i,j)$ by $\operatorname{TB}_{i,j}(X)$. By $\operatorname{TB}(X)$ we mean the truth value of the conjecture for a given $X$ and all $i,j\in{\mathbb{Z}}$. When $X$ is of pure dimension $d$, we define $\operatorname{TB}^{i,j}(X)=\operatorname{TB}_{2d-i,d-j}(X)$ (so the notation is compatible with Poincaré duality for a smooth $X$). The conjecture $\operatorname{TB}$, taken for all smooth projective varieties over ${\mathbb{F}}_{q}$, simultaneously generalizes the Tate–Beilinson and Parshin conjectures [Kahn2005, Conjecture 51 and Definition 55]. See loc. cit. for an overview of the general consequences of these conjectures.
We now give some sufficient conditions for $\operatorname{TB}_{i,j}(X)$ (and related statements) to hold.
Proposition 4.3 ([Broe2025, Proof of Lemma 4.4]).
For $X$ a finite-type, separated scheme over ${\mathbb{F}}_{q}$, $Z\subseteq X$ a closed subscheme, U=XZU=X-Z the open complement, and $j\in{\mathbb{Z}}$, there is a commutative diagram with exact rows
[…]
Here the superscript $\operatorname{Frob}_{q}\sim 1$ denotes the generalized eigenspace of the eigenvalue 1 of $\operatorname{Frob}_{q}$. In particular, $\operatorname{TB}_{2j,j}(Z)$ and $\operatorname{TB}_{2j,j}(X)$ together imply $\operatorname{TB}_{2j,j}(U)$.
Proposition 4.4.
Let $X$ be projective over ${\mathbb{F}}_{q}$, and let $Z\subseteq X$ be a closed subscheme, with U=XZU=X-Z smooth, affine, and of pure dimension $d$.
  1. (i)
    The weak Lefschetz theorem holds for étale homology: for any $j\in{\mathbb{Z}}$, the natural map
    $H_{i}(Z_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(j))\to H_{i}(X_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(j))$
    is an isomorphism for i<d1i<d-1, and surjective for i=d1i=d-1.
  2. (ii)
    In particular, if 2j<d12j<d-1 and $\dim_{{\mathbb{Q}}_{\ell}}H_{2j}(X_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(j))^{\operatorname{Frob}_{q}\sim 1}=1$, then the cycle class map
    $\operatorname{CH}_{j}(Z)\otimes_{{\mathbb{Q}}}{\mathbb{Q}}_{\ell}\to H_{2j}(Z_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(j))^{\operatorname{Frob}_{q}\sim 1}$
    is surjective.
Proof.
[…]
$H_{i}(U_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(j))\cong H^{2d-i}(U_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(d-j))=0$
[…]
Proposition 4.5.
Let $X$ be quasi-projective over ${\mathbb{F}}_{q}$.
  1. (i)
    If $Z$ is a closed subscheme of $X$ with open complement $U$, and $\operatorname{TB}$ holds for any two of $X$, $Z$ and $U$, then it also holds for the third one [Jannsen1990, Theorem 12.7 b)].
  2. (ii)
    If $X$ is smooth, $L/{\mathbb{F}}_{q}$ is a finite extension, and $\operatorname{TB}_{i,j}(X_{L})$ holds (where $X_{L}$ is considered as a scheme over $L$), then $\operatorname{TB}_{i,j}(X)$ holds.
  3. (iii)
    If $X$ is a smooth projective variety with finite-dimensional motive, the Tate conjecture holds for $X$, and for all $i$, $H^{2j}(X_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(j))$ is 1-semisimple, then $\operatorname{TB}(X)$ holds ([Jannsen2007, Theorem 4.12]).
  4. (iv)
    If $X$ is a smooth projective variety, rational and numerical equivalence coincide on $\operatorname{CH}^{j}(X)$, and the Tate conjecture holds for cycles of codimension $j$ in $X$, then $\operatorname{TB}^{2j,j}(X)$ holds [Tate1994, Theorem 2.9].
Proof.
[…]
$H_{i}^{\mathcal{M}}(X,{\mathbb{Q}}(j))\otimes_{\mathbb{Q}}{\mathbb{Q}}_{\ell}$$H_{i}(X_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(j))^{G_{{\mathbb{F}}_{q}}}$$H_{i}^{\mathcal{M}}(X_{L},{\mathbb{Q}}(j))\otimes_{\mathbb{Q}}{\mathbb{Q}}_{\ell}$$H_{i}(X_{\overline{L}},{\mathbb{Q}}_{\ell}(j))^{G_{L}}.$$\scriptstyle{f^{*}}$$\scriptstyle{f^{*}}$
The left arrow is injective by the projection formula (proposition 4.1(iv)), and the bottom arrow is injective by $\operatorname{TB}_{i,j}(X_{L})$, so the the top arrow is injective.
For the surjectivity, consider
$H_{i}^{\mathcal{M}}(X,{\mathbb{Q}}(j))\otimes_{\mathbb{Q}}{\mathbb{Q}}_{\ell}$$H_{i}(X_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(j))^{G_{{\mathbb{F}}_{q}}}$$H_{i}^{\mathcal{M}}(X_{L},{\mathbb{Q}}(j))\otimes_{\mathbb{Q}}{\mathbb{Q}}_{\ell}$$H_{i}(X_{\overline{L}},{\mathbb{Q}}_{\ell}(j))^{G_{L}}.$$\scriptstyle{f_{*}}$$\scriptstyle{f_{*}}$
The bottom arrow is surjective by $\operatorname{TB}_{i,j}(X_{L})$, and the right arrow is surjective by the projection formula, so the top arrow is surjective. ∎
We remark that when $X$ is a smooth projective variety over ${\mathbb{F}}_{q}$, rational and numerical equivalence coincide on $\operatorname{CH}^{j}(X)$ when $j=1$ (by the theory of the Picard variety and [Broe2025, Lemma 2.3]) or when $j=\dim X$ (by the theory of the Albanese variety and proposition 2.4).
Corollary 4.6.
The conjecture $\operatorname{TB}(X)$ holds for any smooth quadric hypersurface $X$ over ${\mathbb{F}}_{q}$.
Proof.
Using proposition 4.5(ii), we may assume that $X$ has a rational point. The claim then follows from [EKM2008, Proposition 70.1] and induction on the dimension of $X$. ∎
Corollary 4.7.
The conjecture $\operatorname{TB}(X)$ holds for a smooth projective variety $X$ over ${\mathbb{F}}_{q}$ in the following cases:
  1. (i)
    $X$ has motive of abelian type and satisfies the Tate conjecture;
  2. (ii)
    $X$ has dimension $\leq 3$, and there exists a dominant rational map $Y\dashrightarrow X$ from a smooth projective variety $Y$, such that $Y$ has motive of abelian type, and satisfies the Tate conjecture for divisors.
Proof.
Claim (i) is clear from proposition 4.5 and the semisimplicity of the $\ell$-adic cohomology of abelian varieties over finite fields, which follows from [Tate1966, page 138] and Chevalley’s theorem [Maculan2018, Theorem 1].
In the situation of (ii), by [Vial2017, Theorem 3.11 and Proposition 3.13], there exists a direct sum decomposition
$\mathfrak{h}(X)\cong Q\oplus R(-1),$
where $Q$ is a direct summand of $\mathfrak{h}(Y)\oplus\mathfrak{h}(C_{1})\oplus...\oplus\mathfrak{h}(C_{m})$ for some curves $C_{1},...,C_{m}$, and $R$ is a direct summand of $\mathfrak{h}(S_{1})\oplus...\oplus\mathfrak{h}(S_{n})$ for some varieties $S_{1},...,S_{n}$ of dimension $\leq 2$ [^3] Note that [Vial2017] uses a “covariant” convention for motives, and the opposite of our sign convention for Tate twists. Also, its definition of “variety” does not require irreducibility.. Theorem 3.12 of loc. cit. additionally implies that the motive of $X$ is finite-dimensional. Using proposition 4.5(ii), we are free to extend scalars and assume that all $C_{i}$ and $S_{j}$ are geometrically connected. It then follows from the decomposition of the motive that $X$ satisfies the Tate conjecture for divisors, and that $H^{2j}(X_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(j))$ is 1-semisimple for $i\leq 1$. The full Tate conjecture for $X$, and the 1-semisimplicity of its cohomology in higher degrees, then follows from the hard Lefschetz theorem [Deligne1980, Théorème (4.1.1)]. We conclude by proposition 4.5(iii). ∎
Corollary 4.8.
The conjecture $\operatorname{TB}(X)$ holds when $X\subset\mathbb{P}^{n+1}$ is a Fermat hypersurface of degree $m$, and $p=\operatorname{char}({\mathbb{F}}_{q})$ satisfies $p^{r}\equiv-1\pmod{m}$ for some $r$.
Proof.
Such an $X$ has motive of abelian type, and satisfies the Tate conjecture [Tate1965, page 102]. The claim thus follows from corollary 4.7. ∎
Corollary 4.9.
The conjecture $\operatorname{TB}(X)$ holds for any smooth cubic hypersurface $X$ over ${\mathbb{F}}_{q}$ of dimension $\leq 3$.
Proof.
The claim is clear for $\dim X=1$, by proposition 4.5. Assume $\dim X>1$. By the same proposition, we may pass to a finite extension and assume that $X$ has a rational point. Then by [Kollar2002, Theorem 1], $X$ is unirational, so corollary 4.7 applies. ∎
Proposition 4.10.
Let $X$ be a smooth cubic hypersurface over $\mathbb{F}_{q}$ of dimension $d$. Then $\operatorname{TB}^{2j,j}(X)$ holds for $j\in\{1,2,d-1,d\}$.
Proof.
By corollary 4.9 we may assume $d\geq 4$. The claim then follows from combining proposition 4.5(iv) with the results of section 2 on Chow groups of cubic hypersurfaces. Specifically, when $j=1$, we use the Tate conjecture for divisors on $X$; when $j=2$, we use the agreement of rational and numerical equivalence on $\operatorname{CH}^{2}(X)$ (proposition 2.2) and the Tate conjecture in codimension 2 for $X$ (proposition 2.7); when j=d1j=d-1 we argue similarly, via corollary 2.5 and proposition 3.3, and when $j=d$ we use the finiteness of the Albanese kernel (proposition 2.4). ∎
We note that Rapoport proved in general that the $\ell$-adic cohomology of a smooth cubic fourfold over ${\mathbb{F}}_{q}$ is a semisimple Galois representation ([Rapoport1972]; see also [Zarhin2004, page 1]).
Using Voevosky’s proof of the Bloch–Kato conjecture [Voe2011], one can get some limited control over the integral motivic cohomology of varieties satisfying the conjecture $\operatorname{TB}$.
Proposition 4.11.
Let $X$ be a smooth projective variety over ${\mathbb{F}}_{q}$ such that $\operatorname{TB}(X)$ holds, and let $i,j\in{\mathbb{Z}}$. Then if either
  1. (i)
    $\dim X\leq 3$, or
  2. (ii)
    $i\leq j+2$,
then $H^{i}_{\mathcal{M}}(X,{\mathbb{Z}}(j))$ is a finitely generated abelian group.
Sketch of proof.
This follows from the argument of Corollaries 69 and 70 of [Kahn2005]. We will fill in some details for the reader who wishes to check them, referring to the statements of loc. cit. as numbered therein. The crucial point is to observe, from the equivalent formulation of the Bloch–Kato conjecture in Lemma 18, that $H^{i}_{\mathcal{M}}(X,\mathbb{Z}(j))$ injects into $H^{i}_{\operatorname{\acute{e}t}}(X,\mathbb{Z}(j))$ for $i\leq j+2$. Here the latter group is étale motivic cohomology. Now use the finite generation of the Weil-étale cohomology of $X$ (proved in Theorem 67) and the long exact sequence of (4.39). Recalling that $H^{i}_{\mathcal{M}}(X,{\mathbb{Q}}(j))\cong H^{i}_{\operatorname{\acute{e}t}}(X,{\mathbb{Q}}(j))$ (Theorem 14), one finds that $H^{i}_{\operatorname{\acute{e}t}}(X,{\mathbb{Z}}(j))$ can only fail to be finitely generated when $i=2j+2$. From this one concludes that $H^{i}_{\mathcal{M}}(X,{\mathbb{Z}}(j))$ is finitely generated whenever $i\leq j+2$ and $j>0$. The lower bound on $j$ is removed through the straightforward observation that when $j\leq 0$, one has
$\operatorname{CH}^{j}(X,n)_{\mathbb{Z}}=\begin{cases}{\mathbb{Z}}&j=n=0\\ 0&\textrm{ otherwise},\end{cases}$
cf. [Bloch1986, page 268]. ∎
4.2. The Beilinson–Bloch conjecture
For $X$ a smooth projective variety over a global field $k$, the $\ell$-adic cohomology groups of $X$ have associated Hasse–Weil $L$-functions. When $\operatorname{char}(k)>0$, these functions are rational by the Grothendieck–Lefschetz trace formula, and when $k$ is a number field, they are conjectured to have meromorphic continuation to $\mathbb{C}$.
Conjecture 4.12 ($\operatorname{BB}^{i}(X)$).
For an integer $i$ with $0\leq i\leq\dim X$,
$\dim_{\mathbb{Q}}\operatorname{CH}^{i}(X)_{0}=\operatorname*{ord}\limits_{s=i}L(H^{2i-1}(X_{\overline{k}},{\mathbb{Q}}_{\ell}),s).$
This conjecture was independently formulated over number fields by Beilinson [Beilinson1987, Conjecture 5.0] and Bloch [Bloch1984, page 94] (and both attributed it to Swinnerton-Dyer, who seems not to have published it). We write $\operatorname{BB}^{i}(X)$ for the truth value of the conjecture for a given $X$ and $i$. Sometimes included along with conjecture 4.12 are the statements that the Chow group with integer coefficients $\operatorname{CH}^{i}(X)_{\mathbb{Z}}$ is finitely generated (the “motivic Bass conjecture”) [^4] Note that Kahn expresses some doubt about the analogous assertion over finite fields in [Kahn2005, Section 4.9.1]., and that the Abel–Jacobi map
$\operatorname{CH}^{i}(X)_{0}\otimes_{{\mathbb{Q}}}{\mathbb{Q}}_{\ell}\to H^{1}_{\operatorname{cont}}(G_{k},H^{2i-1}(X_{\overline{k}},{\mathbb{Q}}_{\ell}(i)))$
is injective [Jannsen1990, Conjecture 9.15].
Little is known in general about the Beilinson–Bloch conjecture, and we will not attempt more than a very limited recollection of existing results (see [LTXZZ2025, Remark 1.3] for some additional background). When $k={\mathbb{Q}}$ or when $\operatorname{char}(k)>0$, the conjecture is known for elliptic curves over $k$ of analytic rank at most one: this claim over $\mathbb{Q}$} follows from [GZ1986], [Koly1988], [Wiles1995], and [BCDT01], and over function fields from [Qiu2022, Theorems 12.1.2 and 12.3.1]. When $k$ is a number field and $i>1$, partial evidence has come from the study of Ceresa cycles (as in [Bloch1984]). The conjecture $\operatorname{TB}$, taken for all varieties over finite fields, is known to imply the Beilinson–Bloch conjecture in positive characteristic ([Jannsen1990, Chapter 12], [Broe2025, Corollary 5.6]), and one can use this approach to prove the conjecture unconditionally in special cases (e.g. in the cases mentioned in the introduction, and in [Ulmer2011, Theorem 3.9.1]). In particular, we have the following criterion, which will be our main tool in proving cases of the Beilinson–Bloch conjecture later in the paper.
Proposition 4.13 ([Broe2025, Remark 5.3 and Corollary 5.6]).
Let $f:X\to C$ be a flat, projective morphism to a smooth, projective, geometrically connected curve $C$ over a finite field ${\mathbb{F}}_{q}$. Assume that the generic fiber $X_{k(C)}$ is a smooth variety over the function field of $C$. Let $Z$ be a finite nonempty set of closed points of $C$, and U=CZU=C-Z, such that $f$ is smooth over $U$. Let $d=\dim X$ and let $0\leq j\leq d-1$. If the cycle class map
$\operatorname{CH}_{d-j}(X_{Z})\to H_{2d-2j}((X_{Z})_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(d-j))$
surjects onto the Galois-invariant subspace, its target $H_{2d-2j}((X_{Z})_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(d-j))$ is 1-semisimple, and $\operatorname{TB}^{2j,j}(X)$ holds, then $BB^{j}(X_{\eta})$ holds, and the Abel–Jacobi map
$\operatorname{CH}^{j}(X_{\eta})_{0}\otimes_{\mathbb{Q}}{\mathbb{Q}}_{\ell}\to H^{1}_{\operatorname{cont}}(G_{k(C)},H^{2j-1}(X_{\overline{k(C)}},{\mathbb{Q}}_{\ell})))$
is bijective. Here $H^{1}_{\operatorname{cont}}$ denotes continuous group cohomology.
We remark that it would follow from the Beilinson–Bloch conjecture that the Albanese kernel of any smooth, projective, geometrically connected variety over a global field is torsion [Beilinson1987, Lemma 5.1].
Tne Beilinson–Bloch conjecture, in the form stated here, is closely related to conjectures of Beilinson, Bloch and Murre describing filtrations on the Chow groups of smooth projective varieties over arbitrary fields. We refer to Jannsen’s article [Jannsen1994] for a wonderful overview of some of the ideas behind and connections between the conjectures.
5. Monodromy of Lefschetz pencils
Here we compile some needed results from [SGA7-II] and [Deligne1980] on the monodromy of Lefschetz pencils, and use them to compute the Hasse–Weil $L$-function arising from the middle cohomology of a Lefschetz pencil over a finite field. This will later shed light on the arithmetic of the intermediate Jacobians of certain cubic threefolds (proposition 7.1).
Let $k$ be a field and $X$ a smooth, projective, geometrically connected variety over $k$ of dimension $d\geq 1$. We fix a closed immersion $X\subset\mathbb{P}^{r}_{k}$ [^5] In concrete examples, the specific closed immersion we consider will typically be clear from context.. Let $\check{\mathbb{P}}^{r}_{k}$ denote the dual projective space parametrizing hyperplanes in $\mathbb{P}^{r}_{k}$, and let $\operatorname{Gr}(1,\check{\mathbb{P}}^{r}_{k})$ denote its Grassmannian of lines. A line $h\in\operatorname{Gr}(1,\check{\mathbb{P}}^{r})(k)$ has an associated axis $A(h)=\bigcap\limits_{t\in\mathbb{P}^{1}(k)}h(t)$, which is a codimension-2 linear subspace in $\mathbb{P}^{r}_{k}$.
Definition 5.1 ([SGA7-II, XVII.2.2]).
A line $h\in\operatorname{Gr}(1,\check{\mathbb{P}}^{r})(\overline{k})$ is called a Lefschetz pencil for $X_{\overline{k}}$ if $h$ satisfies the following conditions:
  1. (i)
    the axis of $h$ intersects $X_{\overline{k}}$ transversely;
  2. (ii)
    there exists a dense open $U\subseteq\mathbb{P}^{1}_{\overline{k}}$ such that for all $t\in U(\overline{k})$, the hyperplane $h(t)$ intersects $X_{\overline{k}}$ transversely;
  3. (iii)
    for all $t\in\mathbb{P}^{1}(\overline{k})-U(\overline{k})$, $h(t)$ intersects $X_{\overline{k}}$ transversely except at a single point, which is an ordinary double point singularity of $X_{\overline{k}}\cap h(t)$ [SGA7-II, Définition XV.1.2.1]. We also refer to such a singularity as a node.
More broadly, we say that $h\in\operatorname{Gr}(1,\check{\mathbb{P}}^{r})(k)$ is a Lefschetz pencil for $X$ if it is a Lefschetz pencil for $X_{\overline{k}}$ in the above sense. We say that $X_{\overline{k}}\subset\mathbb{P}^{r}_{\overline{k}}$ is a Lefschetz embedding if $\operatorname{Gr}(1,\check{\mathbb{P}}^{r}_{\overline{k}})$ contains an open dense subvariety whose closed points consist of all Lefschetz pencils for $X_{\overline{k}}$.
Remark 5.2.
An ordinary double point singularity of a variety $Z$ over $k$ is non-degenerate in the sense of [SGA7-II, Définition XV.1.2.2] iff either $\dim Z$ is odd, or $\operatorname{char}(k)\neq 2$. Several of the results we will state have similar hypotheses on dimension and characteristic, which are partly explained by the need to address the edge case of degenerate ordinary double points.
We proceed to recall a criterion for the existence of Lefschetz pencils (over $\overline{k}$). Let $N$ be the normal bundle of $X$ in the ambient projective space. Then the associated projective bundle $\mathbb{P}(N)$ embeds into $X\times\check{\mathbb{P}}^{r}_{k}$, and its image consists of pairs $(x,H)$, with $H$ a hyperplane tangent to $X$ at $x$. The image of the projection $\phi:\mathbb{P}(N)\to\check{\mathbb{P}}^{r}_{k}$ is called the dual variety of $X$, denoted by $\check{X}$ [SGA7-II, XVII.3.1.3].
Lemma 5.3 ([SGA7-II, Corollaire XVII.3.5.0]).
Assume that $d$ is even, or that $\operatorname{char}(k)\neq 2$. Then $X_{\overline{k}}\subset\mathbb{P}^{r}_{\overline{k}}$ is a Lefschetz embedding if and only if either $\dim\check{X}<r-2$, or the extension of function fields $k(\mathbb{P}(N))/k(\check{X})$ is finite separable.
When $X$ is a smooth cubic fourfold and $\operatorname{char}(k)\notin\{2,3\}$, the conditions of the lemma are satisfied [CP2019, Lemme 3.2].
Fix a Lefschetz pencil $h\in\operatorname{Gr}(1,\check{\mathbb{P}}^{r})(k)$ for $X$, and let $\Delta=\Delta(h)=X\cap A(h)$ be the base locus of the pencil. The blowup $Y$ of $X$ along $\Delta$ is smooth, and admits a morphism $f:Y\to\mathbb{P}^{1}$ such that $f^{-1}(t)\cong X_{\overline{k}}\cap h(t)$ for all $t\in\mathbb{P}^{1}(\overline{k})$ [SGA7-II, XVIII.3.1]. We will slightly abuse terminology and also refer to $f$ as a Lefschetz pencil for $X$. Let $\eta$ be the generic point of $\mathbb{P}^{1}_{k}$, and $\overline{\eta}\to\eta$ a geometric generic point. Let $G_{\eta}=\operatorname{Gal}(\overline{\eta}/\eta)$, and let $V=H^{d-1}(Y_{\overline{\eta}},{\mathbb{Q}}_{\ell})$.
Proposition 5.4.
Assume that one of the following conditions holds:
  1. (i)
    $d$ is odd;
  2. (ii)
    $d$ is even, and the pullback map $H^{d-1}(X_{\overline{k}},{\mathbb{Q}}_{\ell})\to V$ is not surjective.
Then the Leray spectral sequence
$E_{2}^{i,j}:H^{i}(\mathbb{P}^{1}_{\overline{k}},R^{j}f_{\overline{k},*}{\mathbb{Q}}_{\ell})\implies H^{i+j}(Y_{\overline{k}},{\mathbb{Q}}_{\ell})$
degenerates at $E_{2}$, and we have canonical isomorphisms
$E_{2}^{0,d-1}$ $\cong H^{d-1}(X_{\overline{k}},{\mathbb{Q}}_{\ell}),$
$E_{2}^{2,d-1}$ $\cong H^{d-1}(X_{\overline{k}},{\mathbb{Q}}_{\ell}(-1)).$
compatible with the $G_{k}$-action. For $j\neq d-1$, we have canonical isomorphisms
$E_{2}^{0,j}$ $\cong H^{j}(Y_{\overline{\eta}},{\mathbb{Q}}_{\ell}),$
$E_{2}^{2,j}$ $\cong H^{j}(Y_{\overline{\eta}},{\mathbb{Q}}_{\ell}(-1)),$
$E_{2}^{i,j}$ $=0\textrm{ for }i\notin\{0,2\},$
compatible with the $G_{\eta}$-action through the quotient $G_{\eta}\to G_{k}$.
Proof.
This follows from [SGA7-II, Théorèmes XVIII.5.6 and XVIII.6.3.3] and the hard Lefchetz theorem [Deligne1980, Théorème 4.1.1]. ∎
We will need some facts about vanishing cycles, which we now review. The reader is advised that this requires a bit of care with basepoints. Let $m=\left\lfloor\frac{d-1}{2}\right\rfloor$. Recall that for two geometric points $\overline{a}$, $\overline{b}$ of a scheme $Z$, a path in $Z$ from $\overline{a}$ to $\overline{b}$ is an isomorphism between the associated set-valued fiber functors on the category of finite étale covers of $Z$. Let $S$ be the set of points $\overline{x}\in\mathbb{P}^{1}_{\overline{k}}$ such that the fiber $Y_{\overline{x}}$ is singular, and $U=\mathbb{P}^{1}_{\overline{k}}-S$. For each $\overline{x}\in S$, fix a geometric generic point $\overline{\eta}_{\overline{x}}$ of the henselization of $\mathbb{P}^{1}_{\overline{k}}$ at $\overline{x}$. For each path $c$ in $U$ from $\overline{\eta}$ to $\overline{\eta}_{\overline{x}}$, there is an associated vanishing cycle $\delta_{c}\in V(m)$, well-defined up to sign. We define
$E_{\overline{x}}=\{\pm\delta_{c}:c\textrm{ a path from }\overline{\eta}\textrm{ to }\overline{\eta}_{\overline{x}}\}\subseteq V(m),$
following [Deligne1980, Définition 4.2.4]. Let
$C=\bigcup_{\overline{x}\in S}E_{\overline{x}}\subseteq V(m),$
and let $\operatorname{Ev}(Y_{\overline{\eta}})\subseteq V$ be the subspace spanned by $\{\lambda\delta:\lambda\in{\mathbb{Q}}_{\ell}(-m),\delta\in C\}$.
Proposition 5.5 ([Deligne1980, Corollaire 4.3.9]).
The pullback map $H^{d-1}(X_{\overline{k}},{\mathbb{Q}}_{\ell})\to V$ induces a direct sum decomposition
$V\cong H^{d-1}(X_{\overline{k}},{\mathbb{Q}}_{\ell})\oplus\operatorname{Ev}(Y_{\overline{\eta}}).$
Let $G_{\infty}$ be the Galois group of $\overline{k(\eta)}$ over $\overline{k}(\eta)$, and note that the $G_{\infty}$-action on $V$ factors through the surjection $G_{\infty}\to\pi_{1}(U,\overline{\eta})$. Let $\rho:G_{\infty}\to\mathcal{G}$ denote the $G_{\infty}$-representation on $\operatorname{Ev}(Y_{\overline{\eta}})$. This lands in the group $\mathcal{G}$ of automorphisms of $\operatorname{Ev}(Y_{\overline{\eta}})$ which preserve the nondegenerate cup product pairing ${\operatorname{Ev}(Y_{\overline{\eta}})\otimes_{{\mathbb{Q}}_{\ell}}\operatorname{Ev}(Y_{\overline{\eta}})}\to{\mathbb{Q}}_{\ell}(-(d-1))$.
Proposition 5.6 ([Deligne1980, Théorème 4.4.1]).
If $d$ is even, the image $\rho(G_{\infty})$ is open in $\mathcal{G}$.
Remark 5.7.
When $d$ is even, the cup product pairing on $\operatorname{Ev}(Y_{\overline{\eta}})$ is alternating, and $\mathcal{G}$ is the associated symplectic group (which is nonabelian if $\operatorname{Ev}(Y_{\overline{\eta}})$ is nonzero).
Proposition 5.8 ([Deligne1980, Corollaire 4.2.8]).
Suppose that $d$ is even, or that $d$ is odd and $\operatorname{char}(k)\neq 2$. Then for all $\delta,\delta^{\prime}\in C$, there exists $g\in G_{\infty}$ such that $g(\delta)\in\{\delta^{\prime},-\delta^{\prime}\}$. In particular, if one vanishing cycle is nonzero, then they all are.
Fix $\overline{x}\in S$, and a path $c$ from $\overline{\eta}$ to $\overline{\eta}_{\overline{x}}$. Let $I_{c}\subset G_{\eta}$ be the associated inertia group. Recall the specialization map $\operatorname{sp}:H^{d-1}(Y_{\overline{x}},{\mathbb{Q}}_{\ell})\to H^{d-1}(Y_{\overline{\eta}_{\overline{x}}},{\mathbb{Q}}_{\ell})$ [stacks-project, Tag 0GJ2], whose target is identified via $c$ with $V$.
Proposition 5.9.
The specialization map is injective, and its image in $V$ is equal to $V^{I_{c}}$. This in turn is equal to the orthogonal complement of the vanishing cycle $\delta_{c}$ with respect to the cup product $V\otimes_{{\mathbb{Q}}_{\ell}}V(m)\to{\mathbb{Q}}_{\ell}(-(d-m-1))$. If $d$ is even or $\operatorname{char}(k)\neq 2$, the action of $I_{c}$ on $V$ factors through the tame quotient of $I_{c}$.
Proof.
The claims are immediate from [SGA7-II, Théorème XV.3.4]. ∎
Suppose now that $k={\mathbb{F}}_{q}$ is finite. Under mild hypotheses, the above results permit us to explicitly compute the Hasse–Weil $L$-function of the $G_{\eta}$-representation $V$ in terms of the zeta function of $Y$. For a Galois representation $W$ of a finite field ${\mathbb{F}}_{a}$ with $a$ elements, let
$\chi(W,s)=\det(1-\operatorname{Frob}_{a}a^{-s}|W),$
where $\operatorname{Frob}_{a}\in G_{{\mathbb{F}}_{a}}$ is the geometric Frobenius. Let $(\mathbb{P}^{1}_{k})^{0}$ denote the set of closed points of $\mathbb{P}^{1}_{k}$. For each $x\in(\mathbb{P}^{1}_{k})^{0}$, fix a choice of inertia and decomposition groups $I_{x}\subset G_{x}\subset G_{\eta}$, and regard $V^{I_{x}}$ as a $G_{k(x)}$-representation in the natural way. Let
$L(V,s)=\prod\limits_{x\in(\mathbb{P}^{1}_{{\mathbb{F}}_{q}})^{0}}\chi(V^{I_{x}},s)$
be the Hasse–Weil $L$-function of $V$, and
$\Lambda(s)=\prod\limits_{x\in(\mathbb{P}^{1}_{{\mathbb{F}}_{q}})^{0}}\chi(H^{d-1}(Y_{\overline{x}},{\mathbb{Q}}_{\ell}),s)$
be the $L$-function of the sheaf $R^{d-1}f_{*}{\mathbb{Q}}_{\ell}$ on $\mathbb{P}^{1}_{k}$.
Proposition 5.10.
For all $x\in(\mathbb{P}^{1}_{k})^{0}$, we have
$\chi(V^{I_{x}},s)=\chi(H^{d-1}(Y_{\overline{x}},{\mathbb{Q}}_{\ell}),s),$
so that $L(V,s)=\Lambda(s)$. If the hypotheses of proposition 5.4 are satisfied, then
$L(V,s)=\frac{\chi(H^{d}(Y_{\overline{k}},{\mathbb{Q}}_{\ell}),s)}{\chi(H^{d-1}(X_{\overline{k}},{\mathbb{Q}}_{\ell}),s)\ \chi(H^{d-1}(X_{\overline{k}},{\mathbb{Q}}_{\ell}),s-1)\ \chi(H^{d}(Y_{\overline{\eta}},{\mathbb{Q}}_{\ell}),s)\ \chi(H^{d-2}(Y_{\overline{\eta}},{\mathbb{Q}}_{\ell}),s-1)}.$
Proof.
For any $x\in(\mathbb{P}^{1}_{k})^{0}$, proposition 5.9 gives a $G_{k(x)}$-equivariant identification
$H^{d-1}(Y_{\overline{x}},\mathbb{Q}_{\ell})\cong H^{d-1}(Y_{\overline{\eta}},{\mathbb{Q}}_{\ell})^{I_{x}},$
so that the Euler factors of $L(V,s)$ and $\Lambda(s)$ at each closed point agree. By the Grothendieck–Lefschetz trace formula [Milne1980, Theorem VI.13.3],
$\Lambda(s)=\frac{\chi(H^{1}(\mathbb{P}^{1}_{\overline{k}},R^{d-1}f_{\overline{k},*}{\mathbb{Q}}_{\ell}),s)}{\chi(H^{0}(\mathbb{P}^{1}_{\overline{k}},R^{d-1}f_{\overline{k},*}{\mathbb{Q}}_{\ell}),s)\chi(H^{2}(\mathbb{P}^{1}_{\overline{k}},R^{d-1}f_{\overline{k},*}{\mathbb{Q}}_{\ell}),s)}.$ (5.1)
Now assume that the hypotheses of proposition 5.4 are satisfied, so that the Leray spectral sequence discussed there degenerates at $E_{2}$. We then have a direct sum decomposition
$H^{d}(Y_{\overline{k}},{\mathbb{Q}}_{\ell})\cong H^{0}(\mathbb{P}^{1}_{\overline{k}},R^{d}f_{\overline{k},*}{\mathbb{Q}}_{\ell})\oplus H^{1}(\mathbb{P}^{1}_{\overline{k}},R^{d-1}f_{\overline{k},*}{\mathbb{Q}}_{\ell})\oplus H^{2}(\mathbb{P}^{1}_{\overline{k}},R^{d-2}f_{\overline{k},*}{\mathbb{Q}}_{\ell}),$
[…]
6. Lefschetz pencils of some special varieties
Let $k={\mathbb{F}}_{q}$, and again let $X\subset\mathbb{P}^{r}_{k}$ be a smooth, projective, geometrically connected variety of dimension $d$. Fix a Lefschetz pencil $h$ for $X$ which is defined over $k$. Again let $\Delta$ be the base locus of $h$, let $f:Y\to\mathbb{P}^{1}_{k}$ be the family of hyperplane sections of $X$ associated with $h$, and let $Y_{\eta}$ be the generic fiber of $f$. Our goal is to use the criterion of proposition 4.13 to prove cases of the Beilinson–Bloch conjecture for $Y_{\eta}$. This requires understanding the Chow groups of $Y_{\eta}$, and those of the singular closed fibers of $f$. We will give two similar arguments along these lines, first when $X$ is a cubic fourfold, then when $X$ is a more general cubic hypersurface. When $X$ is a threefold with $\operatorname{CH}^{2}(X)$ one-dimensional, we prove that the group of zero-cycles of degree zero of $Y_{\eta}$ is torsion through more elementary means.
Before treating more concrete situations, we include the following general sufficient condition for the Beilinson–Bloch conjecture for $Y_{\eta}$ in high codimension.
Proposition 6.1.
Fix $j$ with $2j>d+1$. Suppose that $\operatorname{TB}^{2j,j}(X)$ and $\operatorname{TB}^{2(j-1),j-1}(\Delta)$ hold. Also assume that
$\dim_{{\mathbb{Q}}_{\ell}}H^{2j}(X_{\overline{k}},{\mathbb{Q}}_{\ell}(j))^{\operatorname{Frob}_{q}^{r}\sim 1}=1$
for all $r>0$, where the superscript $\operatorname{Frob}_{q}^{r}\sim 1$ denotes the 1-generalized eigenspace of $\operatorname{Frob}_{q}^{r}$. Then $\operatorname{BB}^{j}(Y_{\eta})$ holds.
Proof.
By the blowup formula for motives (proposition 3.6), $\operatorname{TB}^{2j,j}(X)$ and $\operatorname{TB}^{2(j-1),j-1}(\Delta)$ together are equivalent to $\operatorname{TB}^{2j,j}(Y)$. Let $Z\subset\mathbb{P}^{1}_{k}$ be a finite nonempty set of closed points, with $U=\mathbb{P}^{1}_{k}-Z$, such that $f_{U}$ is smooth. Using proposition 4.13, it suffices to show that the cycle class map
$\operatorname{CH}_{d-j}(X_{Z})\to H_{2d-2j}((X_{Z})_{\overline{k}},{\mathbb{Q}}_{\ell}(d-j))^{\operatorname{Frob}_{q}\sim 1}$
is surjective. The argument of proposition 4.5 shows that it even suffices to prove the analogous statement after passing to a finite extension $L={\mathbb{F}}_{q^{r}}$, i.e. that
$\operatorname{CH}_{d-j}((X_{Z})_{L})\to H_{2d-2j}((X_{Z})_{\overline{k}},{\mathbb{Q}}_{\ell}(d-j))^{\operatorname{Frob}_{q}^{r}\sim 1}$
is surjective. Fix such an $L$ so that the points of $Z$ are $L$-rational. Then $(X_{Z})_{L}$ is a finite disjoint union of hyperplane sections of $X_{L}$, and the desired surjectivity follows from the weak Lefschetz theorem (proposition 4.4). ∎
Remark 6.2.
For example, from proposition 2.4 and corollary 4.7, one sees that the hypotheses of proposition 6.1 are satisfied when $d\geq 4$, j=d1j=d-1, and $X$ is a Fermat hypersurface which satisfies the Tate conjecture.
Now we turn our focus to smooth cubic hypersurfaces. The following lemma characterizes the singular fibers of their Lefschetz pencils.
Lemma 6.3 ([Huybrechts2023, Corollary 1.5.16]).
If $Z$ is a cubic hypersurface over a field $L$ such that $Z$ is smooth away from a single $L$-rational ordinary double point $z$, then the blowup of $Z$ at $z$ is isomorphic to the blowup of $\mathbb{P}^{\dim Z}_{L}$ along a smooth complete intersection of multidegree (3, 2).
Theorem 6.4.
Assume that $X$ is a smooth cubic fourfold.
  1. (i)
    The conjecture $\operatorname{TB}$ holds for all closed fibers of $f$.
  2. (ii)
    For all $j$, $\operatorname{TB}^{2j,j}(Y)$ holds, and the Beilinson–Bloch conjecture holds for codimension-2 cycles on $Y_{\eta}$. If $\operatorname{TB}(X)$ holds, then $\operatorname{TB}(Y)$ does as well.
Proof.
Let $F$ be a fiber of $f$ over a closed point $t_{0}\in\mathbb{P}^{1}_{k}$, $T$ the regular locus of $F$, and $\operatorname{Sing}(F)=F-T$ the singular locus. As $\operatorname{Sing}(F)$ is either empty or zero-dimensional, $\operatorname{TB}(\operatorname{Sing}(F))$ holds. Thus, by the “two-out-of-three” part of proposition 4.5, $\operatorname{TB}(T)$ is equivalent to $\operatorname{TB}(F)$. By proposition 4.5(ii), to prove $\operatorname{TB}(T)$, it suffices to prove $\operatorname{TB}(T_{M})$ for some finite extension $M/{\mathbb{F}}_{q}$. To prove $\operatorname{TB}(F)$, we therefore can and do assume that $t_{0}$ is ${\mathbb{F}}_{q}$-rational, so $F$ is a cubic threefold over ${\mathbb{F}}_{q}$. If $F$ is smooth, then $\operatorname{TB}(F)$ holds by corollary 4.9, so we are left to consider the case where $\operatorname{Sing}(F)$ consists of one $k$-rational ordinary double point.
Let $\alpha:B\to F$ be the blowup of $F$ along its singular point. By lemma 6.3, $B$ is rational, so $\operatorname{TB}(B)$ holds by corollary 4.7. The exceptional divisor $D$ of $\alpha$ is a smooth quadric surface, so $\operatorname{TB}(D)$ holds by corollary 4.6. Thus $\operatorname{TB}(F)$ holds by the two-out-of-three property.
It remains to prove (ii). We noted in proposition 4.10 that $\operatorname{TB}^{2j,j}(X)$ holds for all $j$. Since $Y$ is the blowup of $X$ along a smooth cubic surface $\Delta$, and $\operatorname{TB}(\Delta)$ holds (corollary 4.9), $\operatorname{TB}^{2j,j}(Y)$ also holds for all $j$. Then $\operatorname{BB}^{2}(Y_{\eta})$ follows from part (i) and proposition 4.13. If $\operatorname{TB}(X)$ holds, we similarly infer $\operatorname{TB}(Y)$. ∎
Theorem 6.5.
Suppose that $X$ is a smooth cubic hypersurface of dimension $d\geq 5$. Then $\operatorname{TB}^{4,2}$ holds for all closed fibers of $f$. If $\operatorname{TB}^{6,3}(X)$ holds (e.g. $X$ is a Fermat cubic hypersurface satisfying the hypotheses of corollary 4.8), then $BB^{3}(Y_{\eta})$ does as well.
Proof.
Let $F$ be a closed fiber of $f$. We want to prove $\operatorname{TB}^{4,2}(F)=\operatorname{TB}_{2d-6,d-3}(F)$. As in the proof of theorem 6.4, by applying proposition 4.3 to the open-closed decomposition of $F$ into its singular and regular locus, and using proposition 4.5(ii), we reduce to the case where $F$ is a cubic hypersurface over ${\mathbb{F}}_{q}$.
If $F$ is smooth, then $\operatorname{TB}^{4,2}(F)$ holds by proposition 4.10. We may thus assume that $F$ is smooth away from a single $k$-rational ordinary double point. With $\alpha:B\to F$ the blowup of the singular point, we see from proposition 3.3, lemma 6.3, and the blowup formula for motives that $\operatorname{TB}^{4,2}(B)$ holds. In detail, we have
$\mathfrak{h}(B)\cong\mathfrak{h}(\mathbb{P}^{d-1})\oplus\mathfrak{h}(S)(-1),$
where $S\subset\mathbb{P}^{d-1}$ is a smooth complete intersection of type (3, 2). Thus $\operatorname{TB}^{2,1}(S)$ implies $\operatorname{TB}^{4,2}(B)$. When $d=5$, it is well-known that $S$ is a K3 surface [Huybrechts2016, Example 1.1.3], so that $\operatorname{TB}^{2,1}(S)$ follows from the known Tate conjecture for $S$ [KM2016, Theorem A.1]. When $d>5$, we observe from the Chow–Künneth decomposition of proposition 3.3 that $\operatorname{TB}^{2,1}(S)$ holds.
The exceptional divisor $D$ of $\alpha$ is a smooth quadric hypersurface, so $\operatorname{TB}(D)$ holds by corollary 4.6. Proposition 4.3 then implies $\operatorname{TB}^{4,2}(B-D)$. Using the diagram of that proposition with respect to the open-closed decomposition $F=\operatorname{Spec}{\mathbb{F}}_{q}\sqcup(B-D)$, we see that $\operatorname{TB}^{4,2}(F)$ holds.
Now assume $\operatorname{TB}^{6,3}(X)$. The blowup formula then shows that $\operatorname{TB}^{6,3}(Y)$ is equivalent to $\operatorname{TB}^{4,2}(\Delta)$, which in turn follows from proposition 4.10. From $\operatorname{TB}^{6,3}(Y)$ and the preceding discussion of closed fibers of $f$, proposition 4.13 yields $\operatorname{BB}^{3}(Y_{\eta})$, as desired. ∎
Remark 6.6.
For $S$ a smooth cubic hypersurface of dimension $\geq 2$ over a global field, there are some cases of the Beilinson–Bloch conjecture for $S$ which are immediate from previously known results and the Betti numbers of $S$. For instance, $\operatorname{BB}^{1}(S)$ follows from the triviality of the Picard variety of $S$. When $d>3$, $\operatorname{BB}^{2}(S)$ follows from proposition 2.2. Likewise, $\operatorname{BB}^{\dim S}(S)$ follows from remark 2.3. Other cases of $\operatorname{BB}^{i}(S)$ are implied by proposition 2.6. In particular, if we take $X$ to be a cubic fivefold over ${\mathbb{F}}_{q}$ which satisfies the hypotheses of theorem 6.5, we find that rational and numerical equivalence agree on the Chow groups (with rational coefficients) of the corresponding cubic fourfold $Y_{\eta}$ over ${\mathbb{F}}_{q}(t)$.
Proposition 6.7.
Suppose that $\dim(X)=3$, and that $\operatorname{CH}^{2}(X)$ is one-dimensional (e.g. $X$ is a smooth Fermat threefold). Then $\operatorname{CH}^{2}(Y_{\eta})$ is one-dimensional. In particular, the group of zero-cycles of degree zero $\operatorname{CH}^{2}(Y_{\eta})_{0,{\mathbb{Z}}}$ is torsion.
Proof.
As $Y$ is the blowup of $X$ along a smooth curve, the blowup formula for motives shows that $\operatorname{CH}^{2}(Y)$ is two-dimensional. Let $D,L\in\operatorname{CH}^{1}(Y)$, such that $D$ is the class of a closed fiber of $f$, and $L$ is ample. Then $D\cdot L$ is numerically nontrivial in $Y$. Since $D\cdot L$ is in the kernel of the surjection $\operatorname{CH}^{2}(Y)\to\operatorname{CH}^{2}(Y_{\eta})$, its target (which is nonzero) must be one-dimensional. ∎
Proposition 6.8.
Assume that $\operatorname{TB}(X)$ holds. Then $\operatorname{CH}^{2}(Y_{\eta})_{\mathbb{Z}}$ is finitely generated.
Proof.
Proposition 4.11 implies that $\operatorname{CH}^{2}(X)_{\mathbb{Z}}$ is finitely generated, and the blowup formula for motives then implies the same for $\operatorname{CH}^{2}(Y)_{\mathbb{Z}}$. The latter surjects onto $\operatorname{CH}^{2}(Y_{\eta})_{\mathbb{Z}}$. ∎
Corollary 6.9.
If $\operatorname{TB}(X)$ holds, and $X$ is either a smooth threefold with $\operatorname{CH}^{2}(X)$ one-dimensional, or a smooth cubic hypersurface of dimension $\neq 4$, then the homologically trivial subgroup of $\operatorname{CH}^{2}(Y_{\eta})_{\mathbb{Z}}$ is finite.
Proof.
This follows from propositions 6.7 and 6.8, and the cases of the Beilinson–Bloch conjecture discussed in remark $6.$ ∎
Remark 6.10.
The statement of proposition 6.8 was previously used in [Kahn2021, Proof of Corollary 1] to prove finiteness of the Albanese kernel of some isotrivial surfaces over global function fields.
Remark 6.11.
Suppose that at least one of the following conditions holds:
  1. (1)
    $X$ is even-dimensional, and the pullback map $H^{d-1}(X_{\overline{k}},{\mathbb{Q}}_{\ell})\to H^{d-1}(Y_{\overline{\eta}},{\mathbb{Q}}_{\ell})$ is not surjective;
  2. (2)
    $X$ is a Fermat hypersurface of degree $m$, where $p=\operatorname{char}(k)\nmid m$ and m1m-1 is not a power of $p$, and the Lefschetz pencil $h$ is “sufficiently generic” (i.e. it lies outside a certain closed subset of $\operatorname{Gr}(1,\check{\mathbb{P}}^{n})$, which depends on $X$).
Then $Y_{\eta}$ is not isotrivial. In the former case, this follows from propositions 5.5 and 5.6. In the latter case, it follows from [OV2007, Example 9]. Taking $X$ to be the Fermat quartic threefold in corollary 6.9, we thus find that for $h$ sufficiently generic, $Y_{\eta}$ is a non-isotrivial quartic K3 surface over ${\mathbb{F}}_{q}(t)$ whose group of zero-cycles of degree zero is finite. As remarked in [Huybrechts2014, Example 9.3], it would be interesting to find an example of a non-isotrivial K3 surface $Z$ over $\overline{{\mathbb{F}}_{q}(t)}$ with $\operatorname{CH}^{2}(Z)_{0}$ trivial.
6.1. Motivic cohomology groups of $Y_{\eta}$
In this subsection, we assume that $X$ is a smooth cubic fourfold over $k={\mathbb{F}}_{q}$ satisfying $\operatorname{TB}(X)$. Using theorem 6.4 and results of Jannsen, we can compute all motivic cohomology groups of $Y_{\eta}$ (with rational coefficients) in terms of certain simpler invariants, whose definitions we now recall.
Let $K=k(\eta)$. For any smooth variety $W$ over $K$ and $b,$ $c\in{\mathbb{Z}}$, define the arithmetic étale cohomology
[…]
Here $\widetilde{W}$ is a choice of smooth proper scheme over a nonempty open $U\subseteq\mathbb{P}^{1}_{k}$ such that $\widetilde{W}\times_{U}K\cong W$, and the colimit is over all nonempty open $U^{\prime}\subseteq U$. As explained in [Jannsen2007, page 20], the arithmetic étale cohomology receives a regulator map
$H^{b}_{\mathcal{M}}(W,{\mathbb{Q}}(c))\to H^{b}_{\operatorname{ar}}(W,{\mathbb{Q}}_{\ell}(c)).$
For $a$, $b$, $c\in{\mathbb{Z}}$, the choice of $\widetilde{W}$ likewise equips $V=H^{b}(W_{\overline{K}},{\mathbb{Q}}_{\ell}(c))$ with an action of the étale fundamental group $\pi_{1}(U,\overline{\eta})$, where $\overline{\eta}$ is a geometric generic point of $U$. We define the arithmetic Galois cohomology of $V$ via
[…]
The arithmetic étale and Galois cohomology groups, and the regulator map, are independent of the choice of $\widetilde{W}$.
We showed that $\operatorname{TB}$ holds for $Y$, and for all closed fibers of $f$. Now [Jannsen2007, Theorem 5.3] yields the below result.
Theorem 6.12.
The following are true.
  1. (i)
    The regulator maps induce isomorphisms
    $H^{i}_{\mathcal{M}}(Y_{\eta},{\mathbb{Q}}(j))\otimes_{\mathbb{Q}}{\mathbb{Q}}_{\ell}\to H^{i}_{\operatorname{ar}}(Y_{\eta},{\mathbb{Q}}_{\ell}(j))$
    for all $i,j\in{\mathbb{Z}}$.
  2. (ii)
    For $i-2j\notin\{0,1\}$, we have $H^{i}_{\mathcal{M}}(Y_{\eta},{\mathbb{Q}}(j))=0$. For $i=2j+1$, there is an isomorphism
    $H^{2j+1}_{\operatorname{ar}}(Y_{\eta},{\mathbb{Q}}_{\ell}(j))\cong H^{1}_{\operatorname{ar}}(G_{\eta},H^{2j+1}(Y_{\overline{K}},{\mathbb{Q}}_{\ell}(j))).$
In particular, Jannsen’s conjecture [Jannsen1990, Conjecture 12.18] holds for $Y_{\eta}$.
We remark that Murre’s conjecture [Murre1993, 1.4], which [Jannsen2007, Theorem 5.3] also implies for $Y_{\eta}$, was previously known for cubic threefolds [dAMS98, Proof of Theorem 4.4].
7. The intermediate Jacobian
Let $X$ be a smooth cubic fourfold over $k={\mathbb{F}}_{q}$, and $f:Y\to\mathbb{P}^{1}_{k}$ a Lefschetz pencil for $X$, with generic fiber $Y_{\eta}$. Then $Y_{\eta}$ is a smooth cubic threefold over $K={\mathbb{F}}_{q}(t)$. Let $A$ be the Albanese variety of the Fano surface of $Y_{\eta}$, and fix a Chow–Künneth decomposition of $Y_{\eta}$ as in proposition 3.3. Passing to cohomology and Chow groups in proposition 3.4 yields
$H^{3}(Y_{\overline{\eta}},{\mathbb{Q}}_{\ell}(2))\cong H^{1}(A_{\overline{\eta}},{\mathbb{Q}}_{\ell}(1))$
and
$\operatorname{CH}^{2}(Y_{\eta})_{0}=\operatorname{CH}^{2}(\mathfrak{h}^{3}(Y_{K}))=\operatorname{CH}^{1}(A)_{0}\cong A(K)\otimes_{\mathbb{Z}}{\mathbb{Q}}.$
It immediately follows that the Beilinson–Bloch conjecture for $Y_{\eta}$ (proved in theorem 6.4) is equivalent to the Birch and Swinnerton–Dyer conjecture for $A$. The latter is in turn equivalent via [KT2003, Theorem on page 541] to the strong Birch and Swinnerton–Dyer conjecture for $A$, which asserts that the Tate–Shafarevich group of $A$ is finite, that the $L$-function of $A$ vanishes at $s=1$ to order $\operatorname{rk}A(K)$, and that the leading Taylor coefficient of the $L$-function satisfies the expected formula.
Let us collect a few other observations about the arithmetic of $A$, which is largely governed by the monodromy of the corresponding Lefschetz pencil $f$.
Proposition 7.1.
The following are true.
  1. (i)
    The $L$-function of $A$ is
    $L(A,s)=\frac{\det(1-\operatorname{Frob}_{q}q^{-s}|H^{4}(Y_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(1)))}{(1-q^{1-s})^{2}}.$
  2. (ii)
    The base change of $A$ to $\overline{K}$ has endomorphism algebra isomorphic to $\mathbb{Q}$, and is not isogenous to an abelian variety defined over $\overline{{\mathbb{F}}}_{q}$.
  3. (iii)
    The conductor of $A$ is the sum in $\operatorname{Div}(\mathbb{P}^{1}_{{\mathbb{F}}_{q}})$ of the points over which $f:Y\to\mathbb{P}^{1}_{{\mathbb{F}}_{q}}$ is not smooth. In particular, $A$ has everywhere semistable reduction.
Proof.
Claim (i) is immediate from proposition 5.10 and the Betti numbers of cubic hypersurfaces. The “big” geometric monodromy of the Tate module of $A$ (propositions 5.5 and 5.6), together with the Tate conjecture for endomorphisms of abelian varieties ([Tate1966], [Faltings1986], [Zarhin1974a], [Zarhin1974b], [MB1985]), yields (ii).
Finally, the conductor of $A$ is computed directly from proposition 5.9, as follows. Fix $t\in\mathbb{P}^{1}_{{\mathbb{F}}_{q}}$ with $Y_{t}$ singular, and $\overline{t}\to t$ a geometric point. Since the computation of the conductor exponent at $t$ is entirely local, we will elide any discussion of paths, and simply fix a vanishing cycle $\delta\in H^{3}(Y_{\overline{\eta}},{\mathbb{Q}}_{\ell}(1))$ for $\overline{t}$. Propositions 5.5 and 5.8 imply that $\delta$ is nonzero. The inertia $I_{\overline{t}}$ acts tamely on the rational $\ell$-adic Tate module $V_{\ell}(A)\cong H^{3}(Y_{\overline{\eta}},{\mathbb{Q}}_{\ell}(1))^{\vee}$ by proposition 5.9, and the inertia-invariant subspace is the orthogonal complement of $\delta$, so the conductor exponent of $V_{\ell}(A)$ at $t$ is
$\dim_{{\mathbb{Q}}_{\ell}}V_{\ell}(A)-\dim_{{\mathbb{Q}}_{\ell}}V_{\ell}(A)^{I_{\overline{t}}}=1.$
The Birch and Swinnerton–Dyer conjecture in positive characteristic is known for isotrivial abelian varieties [Milne1968, Theorem 3] and for abelian varieties of analytic rank zero (for Jacobians over global function fields, combining [Ulmer2011, Formula (2.9.3)] with the arguments of [Tate1966b, Section 4] shows that analytic rank upper bounds algebraic rank; the extension of this inequality to general abelian varieties is well-known folklore). The above proposition thus indicates that the conjecture for $A$ is “hard”, in that it is not implied by either of these criteria. This may be unsurprising, since we used the highly nontrivial Tate conjecture for cubic fourfolds in order to prove it.
Example 7.2.
We explicitly compute the $L$-function of $A$ when $X$ is the Fermat cubic fourfold. Let $\Delta$ be the base locus of $f$, and for simplicity assume that $\Delta$ is isomorphic to a blowup of $\mathbb{P}^{2}$ at six points (which can be guaranteed after passing to a finite extension of ${\mathbb{F}}_{q}$). Using the blowup formula for motives and proposition 7.1, the $L$-function of $A$ can now be calculated directly from the zeta function of $X$. The zeta functions of Fermat varieties were expressed by Weil in terms of Jacobi sums [KS79, Section 2.3]. In particular, when $p\equiv 2$ modulo 3, $H^{4}(X_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(2))$ is generated by algebraic cycles (loc. cit., Theorem 2.10), so that
$\det(1-\operatorname{Frob}_{q}q^{-s}|H^{4}(X_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(1)))$ $=(1-q^{1-s})^{23},$
$\det(1-\operatorname{Frob}_{q}q^{-s}|H^{4}(Y_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(1)))$ $=(1-q^{1-s})^{30},$
$L(A,s)$ $=(1-q^{1-s})^{28}.$
In this case, the Taylor expansion of the $L$-function at $s=1$ is
$L(A,s)=(\ln(q)(s-1))^{28}+O((s-1)^{29}),$
and the Mordell–Weil rank of $A$ is 28. The leading Taylor coefficient suggests that the BSD invariants of this $A$ are as simple as possible.
Remark 7.3.
Combining remark 3.5 with theorem 6.5, one may be able to prove analogues of the results of this section for abelian varieties which appear in the motives of certain cubic fivefolds over ${\mathbb{F}}_{q}(t)$.
Remark 7.4.
We conclude by proving a case of the Tate conjecture. Suppose that $\operatorname{char}(k)\neq 2$. Let $F$ be the Fano surface of $Y_{\eta}$, let $\mathcal{F}\to\mathbb{P}^{1}_{k}$ be a flat projective morphism with generic fiber $F$, and let $V\to\mathcal{F}$ be a projective resolution of singularities of $\mathcal{F}$ (which exists by [CP2008, Theorem 2.1]). The Tate conjecture holds for $F$ [Roulleau2013, Theorem 1], and the Birch and Swinnerton-Dyer conjecture holds for $\operatorname{Alb}(F)$, so the Tate conjecture holds for divisors on $V$ by [Geisser2021, Theorem 1.1]. The hard Lefschetz theorem [Deligne1980, Théorème 4.1.1] in turn implies the full Tate conjecture for $V$.